OpenStudy (anonymous):

Which statement is true about the equation: 3/4z-1/4z+1=2/4z+1 It has no solution. It has one solution. It has two solutions. It has infinitely many solutions.

OpenStudy (anonymous):

@GretaKnows

OpenStudy (anonymous):

this is the same as the other one we did. the equation is equal on both sides 2/4z+1=2/4z+1. so what do you think?

OpenStudy (jackofalltradez):

3/4z - 1/4z + 1 = 2/4z + 1 2/4z +1 = 2/4z + 1

OpenStudy (anonymous):

it has one solution? @GretaKnows

OpenStudy (jackofalltradez):

@kool_22 Nope. It has infinitely many solutions, because z = z

OpenStudy (anonymous):

oh I get it now thanks I gave you a medal and do you want a fan?

OpenStudy (anonymous):

@JackofallTradez